# Boosting Generalization in Diffusion-Based Neural Combinatorial Solver via Inference Time Adaptation

### v2

@@ 4.2 Recursive Denoising-decoding Travel Strategy
Algorithm 1 outlines our proposed 2-phase Inference Time Adaptation framework, which combines recursive denoising-decoding travel strategy with energy-guided sampling. This approach offers a practical and efficient method for achieving zero-shot cross-problem transfer using pre-trained diffusion models, without any additional training costs. Furthermore, the framework exhibits flexibility by incorporating advanced heuristic decoding strategies, such as 2-opt heuristics [39] or Monte Carlo Tree Search (MCTS) [24, 6], to enhance solution quality.
Algorithm 1 Inference Time adaptation for Cross-problem Transfer
@@
1:  $p_{\theta}$: Pre-trained diffusion model
2:  $G^{\prime}$: Target problem instance
3:  $T$: Number of diffusion steps
4:  $i,j$: Number of intermediate steps
5:  $\tau$: Energy guidance temperature
6:  $K$: Number of recursive travel
6:  Optimal solution $\mathbf{y}_{K}$ for problem instance $G^{\prime}$
7:  Initialize $\mathbf{x}_{T}$ with random binary values
8:  for $k=1$ to $K$ do
9:     Re-noise previous estimated solution $\mathbf{y}_{k-1}$ to $\mathbf{x}_{j}$
10:     for $t=j$ to $i$ do
11:        Compute $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ from pre-trained model
12:        Compute energy gradient: $\nabla_{\mathbf{x}_{t}}\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})$
13:        Compute $p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})\propto\exp(-\nabla_{\mathbf{x}_{t}}\phi/\tau)$
14:        Compute posterior $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y^{*}_{G^{\prime}},G^{\prime})$ by Eq (14)
15:        Update next state with Bernoulli Sampling: $\mathbf{x}_{t-1}\sim\text{Cat}(\mathbf{x}_{t-1};p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y^{*}_{G^{\prime}},G^{\prime}))$
@@
17:     Estimate $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{i})=\mathbb{E}_{\widetilde{\mathbf{x}}_{0}\sim p_{\theta}(\widetilde{\mathbf{x}}_{0}|\mathbf{x}_{i})}[\widetilde{\mathbf{x}}_{0}]$
18:     Decode $\widetilde{\mathbf{x}}_{0}$ to the optimal solution $\mathbf{y}_{k}$
19:  end for
20:  return  $\mathbf{y}_{K}$
4.3 Theoretical Analysis of Cross-problem Transfer
Consider the scenario where we aim to transfer knowledge from the original Traveling Salesman Problem (TSP) to its variants: the Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP). The diffusion model, trained on optimal TSP instances, estimates $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ such that, under the assumption of perfect optimization, $p_{\theta}(\mathbf{x}_{0}|G^{\prime})$ approximates $q_{\textup{TSP}}(\mathbf{x}_{0}|G^{\prime})$, where the latter represents the distribution of ground-truth TSP solutions for instance $G^{\prime}$. Through the Bayesian equation established in (11), we expect such a pre-trained prior would substantially enhance posterior sampling solution quality, as TSP shares fundamental structural similarities with PCTSP and OP, thereby encoding relevant domain knowledge. We formalize such similarities in the following analysis. The complete proof is provided in Appendix B. We make use of the following definition to facilitate the analysis.
Definition 4.3 (Marginal Decrease).
$\Delta(S)=\textup{TSP}(V)-\textup{TSP}(V\setminus S).$
Theorem 4.4.
For a non-empty subset of nodes $S\subseteq V$, let $\textup{TSP}(S)$ and $\textup{argTSP}(S)$ denote the optimal cost and optimal tours of TSP on the subgraph specified by $S$. Under Assumptions [^1] We assume simpler setups for PCTSP and OP for clearer presentation of the underlying similarities. These assumptions can be relaxed as discussed in Appendix C. B.2 and B.3, the optimal tours of PCTSP are $\textup{argTSP}(V\setminus S_{\textup{PCTSP}})$, where
$S_{\textup{OP}}\in$ $\arg\min_{S\subseteq V}{\Delta(S)}$
[…]
Theorem 4.4 shows that the optimal solutions of PCTSP and OP are indeed the optimal tours of TSP on some subgraph of $G^{\prime}$, revealing their fundamental structural similarities. We expect that the pre-trained diffusion model is powerful enough to generate high-quality TSP solutions given test instances with various graph sizes. We can therefore understand the inference time adaptation in this scenario as forcing the pre-trained model to focus on a subgraph of $G^{\prime}$. Through this theoretical analysis, we deepen our understanding of how the proposed training-free inference time adaptation works when transferring from a pre-trained network to its variants. Although we focus on PCTSP and OP in this paper, we show that our methodology and insight also hold for more challenging TSP variants such as TSP with time windows in Appendix D.
@@ 5 Numerical Results
Evaluation. Following [10], we generate 1000 test instances for each problem scale: 20, 50, and 100, which denotes the node counts for PCTSP and OP. We evaluate model performance using two primary metrics: average solution cost and optimality gap relative to the exact solution solvers. Additionally, we measure computational efficiency through total training time and per-instance inference time.
@@
Experimental Setup. The proposed Inference Time Adaptation for Diffusion-based solvers (DIFU-Ada) builds upon DIFUSCO’s [6] TSP-trained checkpoints in three different scales without additional training. The backbone archetechture used in diffusion model is the same as DIFUSCO, a 12-layer Anisotropic GNN with a width of 256 hidden nodes. The results of DIFU-Ada are recorded under 100 iterations, with 5-step re-noising and 1-step guided denoising. The decoding strategy used in experiments is greedy decoding. The energy functions used in guided sampling are stated in Proposition 4.1 and 4.2. The guidance temperature $\lambda$ is fixed to be 0.1 and the constraint coefficient $\mu$ is set to be 1 for all problem instances. All experiments are conducted on a single Tesla V100 GPU.
@@
5.1 Cross-problem and Cross-scale Generalization Results
@@
Figure 4: Impact of the number of recursive travel steps in the inference time adaptation on the trade-off between optimality gap (%) and inference time (s) for PCTSP and OP.
@@ 5.2 Ablation Studies
The DIFU-Ada framework integrates two key components: energy-guided sampling and recursive denoising-decoding travel strategy. To assess the contribution of each, we conducted ablation studies. Figure 4 quantitatively illustrates the impact of the recursive denoising-decoding strategy, demonstrating a progressive enhancement in cross-problem transfer capabilities as the number of recursive iterations increases. Figure 5 presents a analysis of performance with or without energy-guided sampling during the recursive iterations. The performance gap observed in these studies show the efficacy of both components and highlight their distinct roles in the overall effectiveness to achieve zero-shot cross-problem transfer.
@@
Figure 5: The Optimality Gap of performing recursive travel strategy with or without energy-guided sampling on PCTSP and OP.
Figure 6: Optimality gap changes with respect to the guided temperature ($-\lambda$ as the x-axis label) on PCTSP and OP.
6 Conclusions
In this work, we introduced an Inference Time Adaptation framework (DIFU-Ada) enabling zero-shot cross-problem transfer for diffusion-based solvers. Through an integration of energy-guided sampling and recursive denoising-decoding travel strategy, our approach successfully transfers pre-trained diffusion models to generate high-quality solutions for TSP variants without additional training. We also provided theoretical analysis in this scenario to deepen our understanding of the problem transfer capability in pre-trained diffusion-based NCO solvers. Numerical experiments on different scales of PCTSP and OP demonstrate that our framework achieves competitive zero-shot performance against existing learning-based methods across various scales.
This work demonstrates the considerable potential of adapting diffusion-based generative NCO solvers to address real-world combinatorial optimization problems, particularly those involving dynamic constraints and objectives. Our approach reduces the computational costs of retraining large-scale neural solvers by providing a flexible, off-the-shelf sampling scheme that enhances both cross-problem and cross-scale generalization capabilities. Nevertheless, several limitations and directions for future research still exist. While our method exhibits promising results on TSP variants, its applicability to a broader spectrum of combinatorial optimization problems beyond routing domains remains to be validated.
@@ Appendix A Problem Descriptions of TSP Variants
Appendix B Proof of Theorem 4.4
To analyze the relation between the optimal solutions of TSP and that of PCTSP/OP, we make use of the following definitions.
Definition B.1 (Marginal Decrease).
@@
Assumption B.2 (PCTSP Setting).
@@
Note that the above assumptions can be relaxed by considering a node-weighted TSP, which can be transformed into a standard TSP as discussed in Appendix C.
@@
[…]
@@
Appendix C Relaxing Assumptions B.2 and B.3
Our analysis in Theorem 4.4 can be extended to the cases with minimum prize constraints or non-uniform rewards by considering node-weighted TSP formulations. The key is that the optimal solutions of PCTSP/OP are the optimal solutions of certain node-weighted TSP over a subgraph as shown below.
@@
Appendix D Adaptation to TSP with Time Windows (TSP-TW)
TSP with Time Windows (TSP-TW) is a variant of TSP that requires each node $i$ to be visited within a given time range $[e_{i},l_{i}]$.
Our approach can also handle TSP-TW through a time-expanded graph transformation:
For each node $i$ with time window $[e_{i},l_{i}]$, we create multiple time-indexed replicas $(i_{e_{i}},i_{e_{i}+1},...,i_{l_{i}})$. The original graph’s edges are then transferred to this expanded representation according to temporal feasibility - if node $i$ connects to node $j$ in the original graph, we connect $i_{t}$ to $j_{t+1}$ in the expanded graph (assuming unit travel time for simplicity).
@@ Appendix D Adaptation to TSP with Time Windows (TSP-TW)
Appendix E Energy Formulations for TSP
In our experimental setup, we employ an anisotropic graph neural network architecture enhanced with edge-gating mechanisms [45, 46]. This encoder transforms the input 2D coordinates and TSP variant paths into a binary edge-selection graph. Additional problem-specific information - such as prizes and penalties in PCTSP and prizes in OP - is incorporated into the gradient calculation of the guidance objective function. The constraint coefficients $\mu$ are chosen to be $1$ in experiments.
Proposition E.1 (Traveling Salesman Problem (TSP)).
Given a complete graph $G=(V,E)$ with edge weights $w:E\rightarrow\mathbb{R}^{+}$, find $\mathbf{x}\in\{0,1\}^{|E|}$ that minimizes:
$\phi(\mathbf{x};G)$ $=f_{\text{cost}}(\mathbf{x};G)-\mu\cdot\log(-g(\mathbf{x};G))$
$\text{where}\quad f_{\text{cost}}(\mathbf{x};G)$ $=\sum_{e\in E}w_{e}x_{e}$
@@ Appendix F Additional Experimental Results
[…]

### v3

@@ 4.3 Recursive Renoising-denoising Travel
Algorithm 1 outlines the framework for Inference-Time Adaptation, combining energy-guided sampling and recursive renoising-denoising travel. It provides a practical and efficient method for zero-shot cross-problem transfer with pre-trained diffusion models at no additional training cost. Furthermore, our framework is model-free and can be augmented with other heuristics, such as 2-opt [35] or MCTS [23, 6], to further enhance final solution quality.
Algorithm 1 Inference Time Adaptation for Cross-problem
@@
[…]
2:$G$: Original problem
[…]
4:$T$: Number of diffusion steps
5:$K$: Number of recursive travel
6:$i$: Number of renoising steps
[…]
7: Optimal solution $y^{(K)}$ for problem instance $G^{\prime}$
8: Initialize $\mathbf{x}^{(0)}_{T}\in\{0,1\}^{N}$
9:for $k=1$ to $K$ do
10:  Re-noise solution $\mathbf{x}_{0}^{(k-1)}$ to some noise level $\mathbf{x}_{i}^{(k)}$
11:  for $t=i$ todo
12:   Compute $p_{\theta}(\mathbf{x}_{t}^{(k)}|G^{\prime})$ from pre-trained model
13:   Compute energy potential: $\nabla_{\mathbf{x}_{t}}\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}^{(k)}_{t});G^{\prime})$
14:   Compute posterior $p_{\theta}(\mathbf{x}^{(k)}_{t-1}|\mathbf{x}^{(k)}_{t},y^{*}_{G^{\prime}},G^{\prime})$
15:   Update next state with Bernoulli Sampling: $\mathbf{x}^{(k)}_{t-1}\sim\text{Cat}(\mathbf{x}^{(k)}_{t-1};p_{\theta}(\mathbf{x}^{(k)}_{t-1}|\mathbf{x}^{(k)}_{t},y^{*}_{G^{\prime}},G^{\prime}))$
@@
17:  Decode $\mathbf{x}^{(k)}_{0}$ to the optimal solution $y^{(k)}$
18:end for
19:return  $y^{(K)}$
@@ 5 Numerical Results
Evaluation. Following [10], we generate 10000 test instances for each problem scale: 20, 50, and 100, which denotes the node counts for PCTSP and OP. We evaluate model performance using two primary metrics: average solution cost and optimality gap relative to the exact solution solvers. Additionally, we measure computational efficiency through total training time and per-instance inference time.
@@
Experimental Setup. The proposed Inference Time Adaptation for Diffusion-based solvers (DIFU-Ada) builds upon DIFUSCO’s [6] TSP-trained checkpoints in three different scales without additional training. The backbone archetechture used in diffusion model is the same as DIFUSCO, a 12-layer Anisotropic GNN with a width of 256 hidden nodes. The results of DIFU-Ada are recorded under 100 iterations, with 5-step re-noising and 1-step guided denoising. The decoding strategy used in experiments is greedy decoding. The energy functions used in guided sampling are stated in Proposition 4.1 and 4.2. The guidance temperature $\lambda$ is fixed to be 0.1 and the constraint coefficient $\mu$ is set to be 1 for all problem instances. All experiments are conducted on a single Tesla V100 GPU. Experimental setup details are provided in Appendix B.
@@
5.1 Cross-problem and Cross-scale Generalization
@@
Figure 3: Ablation studies of the number of recursive travel steps on the trade-off between optimality gap (%) and inference time (s) for PCTSP and OP.
@@
Figure 4: The Optimality Gap of performing recursive travel with or without energy-guided sampling on PCTSP and OP.
@@ 5.2 Ablation Studies
Figure 5: Optimality gap changes with respect to the guided temperature ($\lambda$ as the x-axis label) on PCTSP and OP.
The DIFU-Ada framework integrates two key components: energy-guided sampling and recursive renoising-denosing travel. To assess the contribution of each, we conducted ablation studies. Figure 3 quantitatively illustrates the impact of the recursive renoising-denosing, demonstrating a progressive enhancement in cross-problem transfer capabilities as the number of recursive iterations increases. Figure 4 presents a analysis of performance with or without energy-guided sampling during the recursive iterations. The performance gap observed in these studies show the efficacy of both components and highlight their distinct roles in the overall effectiveness to achieve zero-shot cross-problem transfer.
@@
6 Conclusion
We introduce DIFU-Ada, an inference-time adaptation framework for zero-shot cross-problem transfer in diffusion-based solvers. By integrating energy-guided sampling with a recursive denoising strategy, our approach adapts pre-trained models to generate high-quality solutions for new TSP variants without retraining. Theoretical analysis supports these transfer capabilities, and experiments on PCTSP and OP confirm competitive zero-shot performance against existing learning-based methods.
This work highlights the potential of adapting diffusion solvers for real-world combinatorial optimization problems with dynamic constraints. Our framework offers a flexible, off-the-shelf sampling scheme that avoids costly retraining while enhancing cross-problem and cross-scale generalization. While promising on TSP variants, a key limitation is its validation on routing problems; future work should therefore test its applicability across a broader spectrum of combinatorial domains.
@@ Appendix A Problem Descriptions of TSP Variants
Appendix B Experimental Setup Details
Dataset. We evaluate our approach on two common TSP variants: the Prize Collecting Traveling Salesman Problem (PCTSP) and the Orienteering Problem (OP). Since our method operates in a zero-shot setting, no training or fine-tuning datasets for PCTSP or OP were used. For evaluation, we generated test instances following the standard procedure from prior work [10, 12]. Specifically, for each problem scale $N$, we randomly generated the 2D coordinates of $N$ nodes within a $(0,1)^{N}$ space. For PCTSP, the depot is the first node, and the maximum tour length is set to $N/2$. For OP, prizes are sampled uniformly from $(0,1)$. The reference solutions for evaluation were generated by the exact solver Gurobi for 20-node instances. For larger 50- and 100-node instances where exact solving is intractable, we used high-performance heuristic solvers: ILS for PCTSP and Compass for OP.
[…]
Baselines. We compare DIFU-Ada against three categories of baselines. For each learning-based method, we clarify its training paradigm to ensure a fair comparison:
  • Exact/Heuristic Solvers: For PCTSP, we use Gurobi and Iterated Local Search (ILS). For OP, we use Gurobi, Compass [39], and Tsili [40].
  • Learning-based (Trained/Fine-tuned): These methods are either trained from scratch or fine-tuned on the target problem (PCTSP or OP). This category includes AM [10], MDAM [38], ASP [11] and AM-FT [12].
  • Learning-based (Zero-shot): These methods, like ours, are pre-trained on TSP and applied directly to the target problems without further training. This category includes the original DIFUSCO [6], and T2T [7]. However, these two methods are specifically optimized for solving TSP problems, our proposed DIFU-Ada offers the only zero-shot cross-problem ability without additional training.
Experimental Setup. The proposed Inference Time Adaptation for Diffusion-based solvers (DIFU-Ada) builds upon DIFUSCO’s [6] TSP-trained checkpoints in three different scales without additional training.
The backbone architecture used in diffusion model is the same as DIFUSCO, building upon a GNN Encoder architecture which, for our experiments, is configured with 12 GNN layers and a hidden dimension of 128 for all feature representations. The core of the model is the Gated Graph Convolutional Network. This layer updates both node and edge features by applying a series of linear transformations followed by a sigmoid gating mechanism, with sum aggregation and Layer Normalization used by default. For input processing, the model first converts raw node coordinates into 128-dimensional features using a PositionEmbeddingSine module. The diffusion timestep t is also embedded into a vector representation and injected into each of the 12 GNN layers, allowing the model to condition its predictions on the noise level. Residual connections are employed around each GNN layer to ensure stable training. For computational efficiency, our implementation operates on sparse graph representations.
The results for DIFU-Ada are recorded over 100 recursive iterations, a value found to provide a strong balance between solution quality and inference time. Each iteration consists of a 5-step re-noising phase followed by a 1-step guided denoising phase. We employ a greedy decoding strategy for final solution construction. The energy functions guiding the sampling are defined in Eq. 4.1 and 4.2. To demonstrate robustness without problem-specific tuning, the guidance temperature $\lambda$ was fixed to 0.1 and the constraint coefficient $\mu$ was set to 1 for all instances. All our experiments were conducted on a single NVIDIA Tesla V100 GPU using PyTorch 2.1. Baseline results are taken from their respective publications, and any direct comparison of inference times should consider potential differences in hardware.
Appendix C Theoretical Analysis of Transferring Knowledge from Pre-trained Model
Consider the scenario where we aim to transfer knowledge from the original Traveling Salesman Problem (TSP) to its variants: the Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP). The diffusion model, trained on optimal TSP instances, estimates $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ such that, under the assumption of perfect optimization, $p_{\theta}(\mathbf{x}_{0}|G^{\prime})$ approximates $q_{\textup{TSP}}(\mathbf{x}_{0}|G^{\prime})$, where the latter represents the distribution of ground-truth TSP solutions for instance $G^{\prime}$. Through the Bayesian equation established in (12), we expect such a pre-trained prior would substantially enhance posterior sampling solution quality, as TSP shares fundamental structural similarities with PCTSP and OP, thereby encoding relevant domain knowledge. We formalize such similarities in the following analysis. We make use of the following definition to facilitate the analysis.
Definition C.1 (Marginal Decrease).
$\Delta(S)=\textup{TSP}(V)-\textup{TSP}(V\setminus S).$
Theorem C.2.
For a non-empty subset of nodes $S\subseteq V$, let $\textup{TSP}(S)$ and $\textup{argTSP}(S)$ denote the optimal cost and optimal tours of TSP on the subgraph specified by $S$. Under Assumptions [^1] We assume simpler setups for PCTSP and OP for clearer presentation of the underlying similarities. These assumptions can be relaxed as discussed in Appendix D. C.4 and C.5, the optimal tours of PCTSP are $\textup{argTSP}(V\setminus S_{\textup{PCTSP}})$, where
$S_{\textup{OP}}\in$ $\arg\min_{S\subseteq V}{\Delta(S)}$
 $\qquad\textup{ s.t. }\Delta(S)\geq\textup{TSP}(V)-D_{\textup{OP}}.$
Proof. To analyze the relation between the optimal solutions of TSP and that of PCTSP/OP, we make use of the following definitions.
Definition C.3 (Marginal Decrease).
@@ Appendix C Theoretical Analysis of Transferring Knowledge from Pre-trained Model
Assumption C.4 (PCTSP Setting).
@@
Note that the above assumptions can be relaxed by considering a node-weighted TSP, which can be transformed into a standard TSP as discussed in Appendix D.
@@
 $\ \ \text{ s.t. }\text{TSP}(V\setminus S)\leq D_{\text{OP}}$
$={}$ $\text{TSP}(V)+\min_{S\subseteq V}{\Delta(S)}$ (18)
 $\qquad\qquad\ \ \ \ \text{ s.t. }\Delta(S)\geq\text{TSP}(V)-D_{\text{OP}},$
@@
Theorem C.2 shows that the optimal solutions of PCTSP and OP are indeed the optimal tours of TSP on some subgraph of $G^{\prime}$, revealing their fundamental structural similarities. We expect that the pre-trained diffusion model is powerful enough to generate high-quality TSP solutions given test instances with various graph sizes. We can therefore understand the inference time adaptation in this scenario as forcing the pre-trained model to focus on a subgraph of $G^{\prime}$. Through this theoretical analysis, we deepen our understanding of how the proposed training-free inference time adaptation works when transferring from a pre-trained network to its variants. Although we focus on PCTSP and OP in this paper, we show that our methodology and insight also hold for more challenging TSP variants such as TSP with time windows in Appendix E.
Appendix D Relaxing Assumptions C.4 and C.5
Our analysis in Theorem C.2 can be extended to the cases with minimum prize constraints or non-uniform rewards by considering node-weighted TSP formulations. The key is that the optimal solutions of PCTSP/OP are the optimal solutions of certain node-weighted TSP over a subgraph as shown below.
@@
Appendix E Adaptation to TSP with Time Windows (TSP-TW)
TSP with Time Windows (TSP-TW) is a variant of TSP that requires each node $i$ to be visited within a given time range $e_{i},l_{i}]$. Our approach can also handle TSP-TW through a time-expanded graph transformation:
For each node $i$ with time window $e_{i},l_{i}]$, we create multiple time-indexed replicas $(i_{e_{i}},i_{e_{i}+1},...,i_{l_{i}})$. The original graph’s edges are then transferred to this expanded representation according to temporal feasibility - if node $i$ connects to node $j$ in the original graph, we connect $i_{t}$ to $j_{t+1}$ in the expanded graph (assuming unit travel time for simplicity).
@@ Appendix F Additional Experimental Results
In this part, we provide several additional experimental results.
Table 4 underscores the scalability of our approach on large-scale PCTSP-500 and PCTSP-1000 problems. While heuristic solvers like OR-Tools are prohibitively slow and other learning-based methods require dedicated training to achieve top performance, our training-free DIFU-Ada remains highly competitive. It achieves solution costs (14.5 and 21.7) that are remarkably close to the state-of-the-art trained model, GLOP-S (14.3 and 19.8), while entirely circumventing the need for problem-specific training or fine-tuning. These results validate DIFU-Ada as a robust and efficient framework that balances high performance with zero training overhead, even at a large scale.
@@
Method $downarrow Cost\downarrow Time\downarrow$ Training-free
@@
Table 5 provides a focused evaluation of leading diffusion-based CO solvers on the task of cross-scale generalization for TSP. This comparison demonstrates a key strength of our method: although DIFU-Ada is designed for the more complex challenge of both cross-problem and cross-scale generalization, its performance on the standalone cross-scale TSP task remains highly competitive. The results show that DIFU-Ada achieves performance comparable to the current state-of-the-art, FastT2T, confirming that its enhanced capabilities do not compromise its effectiveness on foundational benchmarks TSP.
@@
Figure 6 illustrates the performance sensitivity of our method to the constraint coefficient $\mu$ for PCTSP-50 and OP-50. The results indicate that while the choice of $\mu$ can slightly alter the convergence speed, the final solution quality remains remarkably stable across a range of values. For both problem types, we observe only a minor performance gap between the different settings. This analysis confirms that our method is robust to the specific choice of $\mu$. Therefore, to simplify the experimental setup and highlight the generalizability of our approach, we use a fixed $\mu=1$ for all experiments, avoiding the need for problem-specific hyperparameter tuning.
@@
Figure 7 provides a visualization of the reverse process for a PCTSP instance. It depicts the evolution of the solution’s adjacency matrix, represented as a heatmap, throughout our recursive renoising-denoising procedure. The process starts from a matrix corresponding to pure uniform noise and progressively denoises it into a structured matrix that represents a feasible PCTSP tour.
Figure 7: Visualized results of applying recursive renoising-denoising to gradually modify the noisy heatmap on PCTSP-20.